40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 8 20:48:06 2017, gautam, Update, SUS, ITMX stuck 
    Reply  Thu Aug 10 09:31:57 2017, Steve, Update, SUS, ITMX sensor voltage 9daysITMX.pngvacGlitchITMX.png
Message ID: 13182     Entry time: Thu Aug 10 09:31:57 2017     In reply to: 13173
Author: Steve 
Type: Update 
Category: SUS 
Subject: ITMX sensor voltage 

There must be some bad connection

Quote:

Somewhere between CDS model restarts and the IFO venting, ITMX got stuck.

I shook it loose using the usual bias slider technique. It appears to be free now, I was able to lock the green beam on a TEM00 mode without touching the green input pointing. The ITMX Oplev spot has also returned to within its MEDM display bounds.

 

Attachment 1: 9daysITMX.png  102 kB  Uploaded Thu Aug 10 10:32:42 2017  | Show | Hide all | Show all
Attachment 2: vacGlitchITMX.png  31 kB  Uploaded Thu Aug 10 11:01:13 2017  | Hide | Hide all | Show all
vacGlitchITMX.png
ELOG V3.1.3-